OpenStudy (anonymous):

x+108+x=170 solve for x

OpenStudy (anonymous):

2x+108=170 2x=62 x=31

OpenStudy (anonymous):

combine like terms 2x+108=170 cancel out 108 2x=170-108 cancel out the 2 x=162/2=31 the answer is 31